A building is expected to require $1,000,000 in capital improvement expenditures in five years (60 months). The building’s net operating cash flow prior to that time is expected to be at least $20,000 at the end of every month. How much of that monthly cash flow must the owners set aside each month in order to have the money available for the capital improvements, assuming monthly interest rate is 1.5%?
Answer #1
Formula :Monthly Installment= P (i/c) / (1 + i/c)^nc - 1
Monthly Installment= 1000000 (0.015) / (1 + 0.015)^60 - 1
Monthly Installment= 15000 / (1.015)^60 - 1
Monthly Installment= 15000 / 2.44321977569 - 1
Monthly Installment= 15000 / 1.44321977569
Monthly Installment to deposit = $10393.43